OpenStudy (anonymous):

one quick question please?????? A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 4. Graph the function using one of the following two options below. One the graph, make sure to label the intercepts.

OpenStudy (anonymous):

a. You may graph your equation by hand on a piece of paper and scan your work. b. You may graph your equation using graphic technology that can be found in the Course Information area.

OpenStudy (ddcamp):

2x + 3y = 1200 3y = 1200 - 2x y = 400 - (2/3)x The y-axis is the number of Tablet cases sold, the x-axis is the number of phone cases sold.